OpenStudy (emilyjones284):

help please..ill give a medal (: Rationalize the denominator and simplify.

OpenStudy (anonymous):

?

OpenStudy (emilyjones284):

\[\sqrt{\frac{ 5m }{ 5n}}\]

OpenStudy (anonymous):

To rationalize, we need to multiply by the denominator, to eliminate the radical on the bottom. First, break this up as follows: \[\frac{ \sqrt{5m} }{ \sqrt{5n} }\] Now multiply top and bottom by rad 5n. This rationalizes it. \[\frac{ \sqrt{5m} }{ \sqrt{5n} }\ \times \frac{ \sqrt{5n} }{ \sqrt{5n} } = \frac{ \sqrt{25mn} }{ 5n } = \frac{ 5\sqrt{nm} }{ 5n } = \frac{ \sqrt{nm} }{ n }\]